Solve 50q-43=52q-81
Answer by edjones(7311)

50q-43=52q-81
81-43=52q-50q add 81-50q to both sides
38=2q
q=38/2 divide both sides by 2
q=19
check
50*19-43=52*19-81
907=907
Ed